Na2CO3(aq) + CaCl2(aq) -------- 2NaCl(aq) + CaCO3 (s)

Calculate the volume (in mL) of 0.100 M CaCl2 needed to produce 1.00g of CaCO (s). There is an excess of Na2CO3

Molar mass of calcium carbonate= 100.09 g/mol

volume of calcium chloride =
1
Expert's answer
2020-11-22T11:26:30-0500

To get the volume of CaCl2 we need to find the number of moles it will have firstly.

So using a mole-mass relationship between CaCl2 and CaCO3 we have;

1mole of CaCl2 produces 100.09g of CaCO3

The the amount of CaCl2 that'll produce 1g of CaCO3= 1×1/100.09=0.01mole.

To get our volume, we'll use the formula; n=CV where; n=number of mole,C=concentration and V=volume in dm^3 or liter.

Making volume subject of the formula we have; V=n/C

V=0.01/0.1=0.1L.

But since the answer is required to be in mL

So we have 0.1×1000=100mL.